please help me fine the point slope form and slope intercept passing through (-2, -5) (6 -5)
so far i have -5 - (-5)/6 - -2 = 0/8 = 0/2 = m
y - (-5) = 0/2 (x - 6)
y + 5 = 0/2 (x - 6)
Is this correct for point slope?
y + 5 = 0/2 (x - 6)
y + 3 - 0/2x +

Answer by checkley71(8403)

(-2,-5)(6,-5) CREATES A LINE WITH A SLOPE OF (-5-(-5))/(6-(-2) OR 0/(6+2) OR 0/8 OR 0
-5=O*-2+M OR -5=M OR WE HAVE Y=0X-5 OR Y=-5 OR A HORIZONTAL LINE PASSING THROUGH THE POINT (0,5)